OpenStudy (anonymous):

Pre-calc help? Check my answer? Find the center, vertices, and foci of the ellipse with equation 3x^2 + 6y^2 = 18 A. Center: (0, 0); Vertices: (0,-sqrt6),(0,sqrt6); Foci: (0,-sqrt3),(0,sqrt3) B. Center: (0, 0); Vertices: (-sqrt6,0),(sqrt6,0); Foci: (-sqrt3,0),(sqrt3,0) C. Center: (0, 0); Vertices: (-6, 0), (6, 0); Foci: (-3*sqrt3,0),(3*sqrt3,0) D. Center: (0, 0); Vertices: (0, -6), (0, 6); Foci: (0,-3*sqrt3),(0,3*sqrt3) I got A as my answer. Am I correct?

OpenStudy (anonymous):

3x^2 + 6y^2 = 18 \[\frac{ 3x^2 }{ 18 }+\frac{ 6y^2 }{ 18 }=1\]\[\frac{ x^2 }{ 6 }+\frac{ y^2 }{ 3 }=1\] so center is (0,0) \[a=\sqrt{6}\]\[b=\sqrt{3}\]

OpenStudy (anonymous):

so the answer is B.

OpenStudy (anonymous):

Oh, alright. Thank you very much for your help!